Does 1/5 and 5/15, make proportions?
Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

No. 5/15 = 1/3, and so is not equal to 1/5.

Answer by mananth(16946)

1/5 * 5/5 = 5/25
so
so it is not in proportion.